# A Distributional Treatment of Real2Sim2Real for Object-Centric Agent Adaptation in Vision-Driven Deformable Linear Object Manipulation

### v2

@@ III Method
5:Run $1$ $\pi_{\boldsymbol{\beta}_{0}}$ rollout in the real env to collect $\mathbf{x}^{r}$
@@
9:     $\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N}\leftarrow$ Run $N$ $\pi_{\boldsymbol{\beta}_{0}}$ rollouts in sim, $\boldsymbol{\theta}\sim p_{i}$
@@
12:     Update reference prior $p_{i}\leftarrow\hat{p}$
13:     $i\leftarrow i+1$
@@
18:Evaluate $\pi_{\boldsymbol{\beta}_{1}}$ in the real env by running $1$ $\pi_{\boldsymbol{\beta}_{1}}$ rollout
@@ III-C Keypoint trajectories in RKHS
The input of the RKHS-net layer is a vector of samples of a distribution. In this work, we compute the distributional embedding of a noisy keypoint trajectory $\mathbf{x}=(\mathbf{x_{1}},...,\mathbf{x_{n}})$, each keypoint defined in a $2$D RGB pixel space.
@@ III-D Policy Learning and Sim2Real Deployment
We use Proximal Policy Optimisation (PPO) [17] as our reference model-free on-policy RL algorithm. PPO maximises the expected reward with a clipped surrogate objective to prevent large updates. It samples actions from a Gaussian distribution, which is parameterised during policy learning. We learn policies in simulation, performing domain randomisation using our LFI reference prior $p$ as a task parameterisation sampler, where $p$ is either the default uniform distribution $\mathit{U}$, or an inferred MoG posterior (Alg. 1, l. 12).
@@ IV-B Simulation & Hardware setup
[…]
The real green target ball is made of plasticine. It is simulated with a fixed rigid sphere of radius $r=1$cm. Since DLO-target physical interactions are irrelevant to our task, this reality gap of the target material is also irrelevant.
@@ IV-D Domain Randomisation setup
Thus, to perform DR for our RL task, we train our policies in vectorised environments [31] and launch parallel instances of our simulation, each with a different $\boldsymbol{\theta}$, sampled by our current reference prior $p(\boldsymbol{\theta})$. To further robustify the Sim2Real transfer of our policies and minimise the need for camera calibration, we introduce a small randomness in our camera and target object positions by uniformly sampling $(x,y,z)$ offsets in the (±$(\pm 0.025,\;\pm 0.025,\;\pm 0.025)$ (m) and (±0.02, 0,±0.02)plus-or-minus0.02 0plus-or-minus0.02($\pm$ 0.02,\;0,\;$\pm$ 0.02)( ± 0.02 , 0 , ± 0.02 ) (m) ranges, respectively.
@@ IV-E Control setup
We command our EEF by sending it cartesian pose commands ($7$D vectors). For our control experiments, we constrain our EEF motion into two dimensions, moving only along the $x$ and $z$ axes by controlling the respective deltas. Thus, our sampled policy actions are $2$D $\langle dx,dz\rangle$ vectors that we sample in the $[-0.06,0.06]$ (m) range to maintain smooth EEF transitions. Our $12$D observations are constructed by flattening and concatenating our EEF’s $\langle x,z\rangle$ $2$D position vector (proprioception) and the $5\times 2$D visual observation, which are the positions of our DLO and target keypoints.
@@
For safety reasons, we restrict our EEF motion within the $\langle x,y,z\rangle\in\langle[0.275,0.6],[-0.1,0.1],[0.1,0.5]\rangle$ (m) world frame coordinates, with the robot arm based on $(0,0,0)$. Whenever the EEF leaves this designated workspace, the episode ends as a failure with a reward of $-1$.
@@ IV-F Parameter Inference & Policy Algorithms training
For parameter inference, we implement the BayesSim-RKHS variant [14] (Fig. 2). Our RKHS-net layer embeds the $5$ observed keypoints using $500$ random Fourier features and cos-only components. Our MDNN models $4$ mixture components with full covariance and uses $3$ fully connected layers of $1024$ units each. We use Adam optimiser with a learning rate of $1e$$-6$. As in Alg. 1, line 8, we approximate our posterior $\hat{p}(\boldsymbol{\theta})$ through $15$ iterations, each increasing the training set with $100$ more trajectories, whose parameters $\boldsymbol{\theta}$ have been sampled using the latest posterior.
@@ V-C Object-centric Agent Performance
[…]
@@
[…]
@@
Figure 5: EEF trajectories during the deployment of $6$ policies trained using different domain distributions (col. 1-6) in real world (rows 1-4; $4$ real DLOs) and in simulation (row 5; median DLO). We repeat each deployment $4$ times and average the measured accumulation of commanded EEF translations along the $x$ and $z$ axes (commanded actions, $\langle dx,dz\rangle$), with shaded regions reporting standard deviation. Colorbars illustrate DLO keypoints pixel-distance to the target $d_{t}$, reversing reward function for $d_{\text{thresh}}=1.5$. The rightmost column shows the averaged rewards per timestep for the respective row’s trajectories.
@@
Figure 7: Dynamic Time Wrapping (DTW) heatmap of real and sim EEF trajectories (accum. commanded actions) over PPO-$\{\mathit{U},\mu,0,1,2,3\}$ evaluations.

### v3

@@ III Method
5:Run 11 $\pi_{\boldsymbol{\beta}_{0}}$ rollout in the real env to collect $\mathbf{x}^{r}$
@@
9:  $\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N}\leftarrow$ Run NN $\pi_{\boldsymbol{\beta}_{0}}$ rollouts in sim, $\boldsymbol{\theta}\sim p_{i}$
@@
12:  $i\leftarrow i+1$
13:  Update reference prior $p_{i}\leftarrow\hat{p}$
@@
18:Evaluate $\pi_{\boldsymbol{\beta}_{1}}$ in the real env by running 11 $\pi_{\boldsymbol{\beta}_{1}}$ rollout
@@ III-C Keypoint trajectories in RKHS
The input of the RKHS-net layer is a vector of samples of a distribution. In this work, we compute the distributional embedding of a noisy keypoint trajectory $\mathbf{x}=(\mathbf{x_{1}},...,\mathbf{x_{n}})$, each keypoint defined in a 22D RGB pixel space.
@@ III-D Policy Learning and Sim2Real Deployment
We use Proximal Policy Optimisation (PPO) [17] as our reference model-free on-policy RL algorithm. PPO maximises the expected reward with a clipped surrogate objective to prevent large updates. It samples actions from a Gaussian distribution, which is parameterised during policy learning. We learn policies in simulation, performing domain randomisation using our LFI reference prior $p$ as a task parameterisation sampler, where $p$ is either the default uniform distribution $\mathit{U}$, or an inferred MoG posterior (Alg. 1, l. 13).
@@ IV-B Simulation & Hardware setup
For our experiments, we manufacture $4$ blue DLOs using Shore hardness A-40 (DLO-0 with len. 200mm), 00-20 (DLO-1 with len. 200mm and DLO-3 with len. 290mm) and 00-50 (DLO-2 with len. 270mm) silicone polymers [33]. Similarly to simulation, our real DLOs are shaped as grids, with negligibly small height and width at $15\text{mm}$ each. Table I summarises our DLO indexes, parameterisations, relative hardness descriptions, and mass. Items are sorted primarily on increasing length and then on increasing softness (as medium soft $\rightarrow$ soft $\rightarrow$ extra soft). We will be referencing DLOs using their index, or as $\langle\text{Length}\rangle;\langle\text{Shore Hardness}\rangle$”.
The real green target ball is made of plasticine. It is simulated with a fixed rigid sphere of radius r=1r=1cm. Since DLO-target physical interactions are irrelevant to our task, this reality gap of the target material is also irrelevant.
@@ IV-D Domain Randomisation setup
[…]
@@ IV-E Control setup
We command our EEF by sending it cartesian pose commands (77D vectors). For our control experiments, we constrain our EEF motion into two dimensions, moving only along the $x$ and $z$ axes by controlling the respective deltas. Thus, our sampled policy actions are 22D $\langle dx,dz\rangle$ vectors that we sample in the [0.06,0.06][-0.06,0.06] (m) range to maintain smooth EEF transitions. Our 1212D observations are constructed by flattening and concatenating our EEF’s $\langle x,z\rangle$ 22D position vector (proprioception) and the $5\times 2$D visual observation, which are the positions of our DLO and target keypoints.
@@
For safety reasons, we restrict our EEF motion within the $\langle x,y,z\rangle\in\langle[0.275,0.6],[-0.1,0.1],[0.1,0.5]\rangle$ (m) world frame coordinates, with the robot arm based on $(0,0,0)$. Whenever the EEF leaves this designated workspace, the episode ends as a failure with a reward of 1-1.
@@ IV-F Parameter Inference & Policy Algorithms training
For parameter inference, we implement the BayesSim-RKHS variant [14] (Fig. 2). Our RKHS-net layer embeds the $5$ observed keypoints using $500$ random Fourier features and cos-only components. Our MDNN models $4$ mixture components with full covariance and uses $3$ fully connected layers of $1024$ units each. We use Adam optimiser with a learning rate of 1e1e6-6. As in Alg. 1, line 8, we approximate our posterior $\hat{p}(\boldsymbol{\theta})$ through $15$ iterations, each increasing the training set with $100$ more trajectories, whose parameters $\boldsymbol{\theta}$ have been sampled using the latest posterior.
@@ V-C Object-centric Agent Performance
We train (Fig. 4) $4$ PPO policies using our $4$ inferred MoG posteriors (Fig. 3) for domain randomisation. We also train a policy by performing DR over a uniform distribution ($\text{PPO-}\mathit{U}$) and a policy which assumes that the simulated DLO is parameterised according to the median of our parameter space ($\text{PPO-}\mu$). We evaluate these $6$ policies in each one of our $4$ real-world DLOs, repeating each evaluation $4$ times for a total of $96$ sim2real policy deployments. Similarly, we also evaluate in simulation over a DLO-$\mu$ parameterised per our parameter space median $\mu$.
@@
We see $\text{MoG}_{\text{DLO-0}}$ policy ($\text{PPO-}0$) showing the tightest “roaming pattern” (1) to maximise the shorter and stiff DLO-0 reward close to the target. $\text{PPO-}0$ also follows the same first half of the trajectory (2) for the very soft DLOs 1 and 3. Similarly, the $\text{MoG}_{\text{DLO-1}}$ policy ($\text{PPO-}1$), follows a very similar first third of the trajectory (3) for the softer DLOs 1, 2 and 3, with some material-related variance, and then attempts a similar “loop pattern” for the last $6$ reaching steps with the softest DLOs 1 and 3 (4). In a related adaptation, we see that although both $\text{PPO-}1$ and $\text{MoG}_{\text{DLO-3}}$ policy ($\text{PPO-}3$) follow a similar pattern for the latter half of the trajectory for the shorter and stiff DLO-0, the PPO-1 pattern looks cleaner, due to being inherently conditioned on shorter DLOs (5).
For the $\text{MoG}_{\text{DLO-2}}$ policy ($\text{PPO-}2$), we see a similar first half of the trajectory for the softest DLOs 1 and 3 (00-20), as well as for DLO-2 (00-50), although with greater variance (6). We also see a similar reaching pattern in the last $6$ steps for the longer DLOs 2 (270mm) and 3 (290mm) (7), with differences attributed to their different softness. PPO-2 and PPO-3 also exhibit a similar reaching pattern in the last $6$ steps for the shorter and softer DLO-1 (8). The most distinct sign of real-world adaptation to domain distribution occurs in PPO-3 trajectories maintaining a greater distance between the EEF and the table than any other DR policy (9). For DLO-$\mu$, reasonably, $\text{PPO-}\mu$ is the most consistent (least variance). Notably, all simulated rollouts require a smaller sum of commanded $\langle dx,dz\rangle$ to approach the target, due to the perfect dynamics of the robot and controller simulation.
Fig. 7 shows the space of the mean EEF trajectories ($4$ rep. average) as a similarity heatmap, to illustrate the variation in the policy space for our $6$ policies evaluated over our $4$ real and $1$ simulated (median) DLOs. $\mathit{U},\mu,0,1,2,3$ denote the respective PPO trajectories. For DLO-0, we see that the PPO-$\{0,1\}$ trajectories, both conditioned on equally short 200mm DLOs, show the greatest similarity. For DLO-1, we see that the $\text{PPO-}\{\mathit{U},2\}$ trajectories are the most similar on average. For the longer DLOs 2 and 3, we have PPO-$\{1,2\}$ being the most similar on average. For DLO-$\mu$ (sim), PPO-$\{1,3\}$ trajectories are the most different, followed by $\text{PPO-}\{\mathit{U},\mu\}$ and $\text{PPO-}\{\mathit{U},1\}$. PPO-$\{1,2\}$ are again the most similar. The greater similarity among all trajectories is due to the median parameters of the DLO, which shows the importance of DLO and domain distribution alignment.
@@
Figure 5: EEF trajectories during the deployment of 66 policies trained using different domain distributions (col. 1-6) in real world (rows 1-4; 44 real DLOs) and in simulation (row 5; median DLO). We repeat each deployment 44 times and average the measured accumulation of commanded EEF translations along the $x$ and $z$ axes (commanded actions, $\langle dx,dz\rangle$), with shaded regions reporting standard deviation. Colorbars illustrate DLO keypoints pixel-distance to the target $d_{t}$, reversing reward function for $d_{\text{thresh}}=1.5$. The rightmost column shows the averaged rewards per timestep for the respective row’s trajectories.
@@
Figure 7: Dynamic Time Warping (DTW) heatmap of real and sim EEF trajectories (accum. commanded actions) over PPO-$\{\mathit{U},\mu,0,1,2,3\}$ evaluations.